# Toward the Categorical Data Map

### v3

@@ IV Interpreting the Categorical Data Map
In the following, we perform a case study on cluster and attribute analysis, using the Property Sales dataset [13] (see Fig. 2) to show how to interpret emerging patterns for cluster, outlier, and similarity analysis. We chose this dataset because of its simplicity and ability to explain concepts. However, it lacks the complexity of large categorical datasets, which we will address in an expert study (see section V).
@@
Cluster Analysis There exist a total of $\Pi_{n\in\{1,\dots,|\mathcal{A}|\}}\,\left|\mathcal{C}_{n}\right|$ possible data items, given that all combinations of attributes are allowed, resulting in an exponential growth in the number of possible and unique data items. Hence, we can assume that there are dependencies and relationships among the categories contained in a dataset impacting their their distribution. This means that groups of subsets that share a set of attributes should form perceivable structures (i.e., clusters) when projected using dimensionality reduction methods. Thus, our approach benefits from and leverages the sparsity of categorical data.
@@
For the Property Sales dataset, there are ten clusters (see Fig. 2 (1)). There is a symmetric split along the center of the projection. Given the size of this dataset, we can observe that the two attributes Purchaser Currently Living In and Location of Property dominate the appearance of the project. Looking at glyph sizes, the categories Private Propriety and Central often occur together while {Private Propriety, Central, Condomimum} is the largest unique subset.
@@
Attribute Analysis: By encoding the attribute values in the background, we enable users to analyze the distribution of subsets in the projection with respect to one or two attributes. For the Property Sales dataset, we found that the attribute Purchaser Currently Living In creates a clear and straight division between subsets (see Fig. 2 (2)). We can also see a second level of grouping by the Location of Property attribute forming a close to orthogonal split in the project, which can be spotted with our visualizations (see Fig. 2 (3) and Fig. 4). We also found that the appearance of the partitioning depends a lot on the selected attributes. When observing the layout enrichment, attributes present themselves on a spectrum from a few clearly separated groups to intermingled and highly fractured appearances. Property Type Purchased does not contribute to the clustering of elements. We can also see that for the Location of Property attribute, the areas of the categories North and West are disjointed, which is reflected in both measures, hinting at a harder-to-interpret layout enrichment and ant-tribute not contributing to cluster cohesion.
Figure 8: Categorical Data Map visualizations of the Titanic dataset [65] using MDS and Overlap coefficient [46]. (1) The visualization shows six clusters and two outliers. The largest cluster is the subset of Adult, Male, Perished (at the top). The background encoding shows that the Survived and Sex attributes are relevant for this dataset, clearly separating the data items. For Sex, the separation is left and right. (2) For Survived, the separation is bottom-left to top-right. (3) The Age dimension also yields a separation, while (4) Class shows no clear structure.
@@
Figure 9: Two visualizations of the Titanic dataset [65]. A split Euler diagram without the Age attribute (left) and an overlap reduced Parallel Sets visualization (right) with very thin ribbons. Both have drawbacks with a small dataset and do not scale well with an increasing number of attributes.
V Evaluation
@@
We compare our Categorical Data Map directly to existing visualization of categorical data, such as Parallel Sets [9, 8]. Additionally, we performed an expert study on the Titanic [65] and Mushroom [69] datasets with five data scientists.
V-A Comparison to Euler Diagrams and Parallel Sets
For categorical data, each data point has exactly one category for each attribute, while in Euler diagrams, the number of sets an element is included is not restricted, i.e., it could be in less. Thus, to truthfully represent categorical data in Euler diagrams, there need to be $\Sigma_{a_{i}\in\mathcal{A}}\,|\mathcal{C}_{i}|$ sets, i.e., one set for each category of all attributes. Euler diagrams may require the selection of specific subsets of attributes and, therefore, are less suitable for exploratory data analysis. For highly intersecting sets, automatic layout methods might not create a single diagram [12]. We show an example of an automatically generated split Euler diagram for the Titanic dataset in Fig. 9 (left). The attribute Age was removed to reduce the diagram’s complexity. The Titanic dataset requires ten sets. However, even with eight sets, the visualization is disjointed.
@@ V-A Comparison to Euler Diagrams and Parallel Sets
Parallel Sets are alternative categorical sets visualization, combining principles from stacked bars and parallel coordinate plots. Fig. 9 (right) shows the Titanic dataset in a through overlap reduced readability improved Parallel Sets visualization. Small subsets are represented as very thin ribbons on the lowest level, which can be hard to perceive. Visualizing the Mushroom dataset with classical Parallel Sets is not visually feasible since it will have 22 ribbon layers and 8123 subsets on the lowest level (see supplementary material). Alsakran et al.  [70] addressed this issue by only visualizing 2-dimensional subsets in a modified Parallel Sets visualization. However, the relation between 2-dimensional subsets is lost. Thus, we argue that Euler diagrams and Parallel Sets, as examples of established visualizations for categorical data, do not scale with increasing dataset size or, more specifically, the number of attributes.
V-B Qualitative Expert User Study
@@
To evaluate the Categorical Data Map we performed a paired analytics study [71]. We conducted an expert study with five data scientists, E1–E5, with varying backgrounds. All participants were Ph.D. candidates and students. All were male, and the age range was 25 to 30 years. All experts had experience in the area of information visualization and visual analytics. During the study, we asked the experts to verbalize their thought process so that we could capture it.
@@
All trials followed a predefined structure and took between 43 and 57 minutes. The study was conducted in German. The study started with an introduction to the Categorical Data Map using the Property Sales dataset by Hassan et al. [72] shown in Fig. 2 and included a description of the square area glyph, layout enrichment, and interactions to introduce the expert to the prototype. After the introduction, the experts had the opportunity to ask questions regarding our approach.
@@
Titanic Dataset: The experts had to analyze the Titanic dataset [65] using the Categorical Data Map shown in Fig. 8. E1–E5 were able to locate the largest subset {Male, Perished, Adult, Crew} by looking at the vitalization without any additional interaction (see Fig. 8 (1)). E1–E5 used Lasso selection to find and validate that the largest subset regarding three attributes is {Male, Perished, Adult} (see Fig. 8 (2)). Additionally, E1–E5 were able to find six clusters and two outliers. E1, E3, and E4 found that the outliers were subsets representing the subsets for the {Survived, Child, Crew} categories (see Fig. 8 (3)). E1, E3, and E5 commented on the high number of perished males and the large number of casualties among the {Male, Crew}. E1–E5 used the layout enrichment to navigate and reason about the location of subsets, including the Class attribute (see Fig. 8 (4)). E2 commented on the close to orthogonal split in the projection between Sex and Survived shown in Fig. 8 (1) and (2).
@@
Figure 10: Categorical Data Map visualizations of the Mushroom dataset [69] using the MDS and Overlap coefficient. (1) Two poisonous mushrooms very similar to edible mushrooms. (2) Comparing class and odor reveals that the poisonous outlier has a pungent odor. Continued analysis reveals that mushrooms with an unpleasant smell are poisonous. (3) After the selection of a cluster, the ring-type is identified as a defining characteristic for the cohesion of visible clusters and is used as a property for the classification of mushrooms. (4) Selecting two poisonous cluster, reveals that the vast amount of poisonous mushrooms are silky at the stalk-surface-below-ring, while there exist very few silky mushrooms that are edible.
@@
Mushroom Dataset: E1–E5 had the opportunity to perform an open exploration task and were only given the information that the dataset is about mushrooms and that the class attribute indicates their poisonousness. The glyph was replaced with a simple black dot. E1–E5 perceived five clusters right at the outset. E2 and E3 used the Lasso selection together with layout enrichment to determine differentiating categories for cluster separation, e.g., evanescent, large, and pendent for the ring-type attribute (see Fig. 10 (3)). E1–E5 found the poisonous outliers nested in the group representing edible mushrooms (see Fig. 10 (1)) being poisonous mushrooms very similar to edible ones. E1–E5 found the general rule that mushrooms with an odor of fishy, foul, musty, spicy, and other unpleasant smells indicate a poisonous mushroom (see Fig. 10 (2)). During the open exploration task, E1, E2, and E3 found the rule without additional information. E4 and E5 needed help to find the class and oder combination. However, E4 and E5 could deduce the rule by only interpreting the plot. E3, being the quickest in exploring the dataset, found that stalk-surface-below-ring is silky for the majority of poisonous mushrooms and that the stalk-surface-below-ring is largely smooth for edible ones. E3 found that poisonous mushrooms generally have a silky stalk-surface-below-ring and few edible ones are silky but rather smooth (see Fig. 10 (4)).
@@
General Comments: Before concluding the study, the participants were asked to comment on their preferences for the available glyph designs. E1, E2, and E4 preferred a circular glyph design (see Fig. 3 (c) and (d)) over a square design. E3 and E5 preferred square glyph designs (see Fig. 3 (a) and (b)). E1 and E3 found that the area-based glyphs are inferior to the alternative designs for reading off precise subset sizes. E1 found the glyph hard to interpret because of multiple possible interpretations if rotated. E1 commented that the layout enrichment is very useful for navigation and orientation and helps to perceive the impact on category groups. However, E1 also noted that the layout enrichment does not indicate the number of data items with the category. E3 mentioned a general preference for the map metaphor by being helpful for orientation among different subsets. E2 mentioned potential scalability issues with the glyph for large datasets, e.g., for a high number of attributes and proposed semantic zoom as a potential option. E1–E5 commented that ordering attributes according to their fracturedness was understandable and useful. During the general questions at the end, E2–E4 freely explored plots created with other distance measures and DR methods. E3 commented that the result of MCA-based plots was hard to interpret, noticing the disjointed layout enrichment. E1 mentioned issues with the encoding of categories, such as the category North not being located north of the plot or the category brown not having the color brown, and suggested being able to change the color of a category manually.
VI Discussion and Future Work
@@
In this section, we discuss lesson-learned findings regarding our approach and reflect on the design decisions, as well as computational complexity and future work.
@@
Visualizing Attributes and Categories: We initially used circular glyphs as shown on Fig. 3 (c) and (d), which had the benefit of using the available space effectively since overlap minimization relative to the radius is straightforward to implement. The subset size encoding by the arc around the circle enables finer-grained distinction of sizes since it offers more space. However, during the design phase, users misinterpreted the circle segments as pie charts, a common method for displaying categorical data. Thus, we decided to circumvent this common misconception by using square-based representation for the categorical subsets. However, three out of five experts preferred a circular glyph design.
@@
Encoding of Subset Sizes: We evaluated four different visual encodings for the size of a categorical data subset (see Fig. 3). The area-based glyph makes it easier to perceive subset sizes at a glance, and thus, a user can spot the distribution of the dataset directly, but it suffers from overlap, especially for tight clusters. There is a benefit of applying methods to reduce the overlap. We are able to largely mitigate the overlap problem with the encoding-based design since it requires less space but the assessment of subset sizes takes more effort. Finally, all glyph designs benefit from a mechanism that moves the currently selected subset to the top, such that all dimensions can be equally well observed. With interaction, it is possible to remove the subset size information altogether. However, this limits analysis to datasets and tasks where the subset sizes it not important, e.g., the Mushroom dataset.
Encoding of Attributes Into the Background: Fig. 8 shows that encoding an attribute into the visualization gives insight into the topology of the projection. We could also show the benefit of encoding multiple dimensions into the background to allow for a more complex representation of the topology.
@@
We found that the number of categories does weakly influence the fracturedness of an attribute. However, the main factor is the number of subsets containing the attribute, i.e., an attribute with two categories and an occurrence and a half and half split among the subsets will yield a low fracturedness. With increased imbalance between the categories, the fracturedness may increase if other more balanced attributes are present.
We also found that a force-directed graph layout will preserve the location of projected points better than Lloyd’s algorithm. Lloyd’s algorithm changes the appearance of clusters in a less predictable way and changes the structure where overlap is not a problem. Thus, we suggest the use of force-directed graph layout over Lloyd’s algorithm.
@@
We discussed the use of weighted Voronoi diagrams [74] to better reflect the subset size in the background encoding. The use of a weighted Voronoi diagram will conflict with the truthful representation of areas assigned to a specific category, more specifically, for imbalanced datasets, the area of one Voronoi cell extends below the point of its neighbors. This behavior makes the layout enrichment hard to interpret since points are placed inside an area representing a category they do not belong to. For datasets with only unique entries, the weight Voronoi diagram will be identical to the regular Voronoi diagram. To organize subsets, we also considered Voronoi Treemaps [75]. However, Voronoi Treemaps require a hierarchical structure, just like regular Tree Maps [76] and, thus, cannot be applied to categorical data without additional information to derive a hierarchy of attributes.
@@
Computational Complexity: Computationally, the number of data records $n$ poses potential limitations. The time complexity of projecting data is determined by the dimensionality reduction methods. However, since categorical data sets are sparse, as discussed in section IV, the number of projected subsets is significantly lower than that of data records. The Voronoi diagram calculation and the corresponding Delaunay triangulation are both in $O(n~{}log(n))$. The time complexity of calculating the fracturedness measures depends on the number of vertices and edges of the Delaunay triangulation, which will have n vertices and $3n-3-h$ edges, where $h$ is the number of vertices on the convex hull. The time complexity of calculating is edge-based fracturedness is based on enumerating all edges of the Delaunay triangulation and has a time complexity of $O(|E|)$. The time complexity of calculating component-based fracturedness is dependent on the algorithm for determining the number of components. We use a depth-first search-based approach with time complexity of $O(|V|+|E|)$. Thus, the dimensionality reduction method employed poses the highest contribution to the time complexity, e.g., for MDS $O(n^{3})$ and for t-SNE $O(n^{2})$, which will be the overall time complexity.
Future Work: One interesting observation we made was that, in clusters, flipping neighboring subsets and their representing cells could decrease fracturedness for an attribute while only marginally increasing the fracturedness of other attributes. Thus, flipping neighboring subsets in clusters can improve the layout for a given attribute. We think that the concept of fracturedness can be transferred to high-dimensional space when analyzing categorical data. Such a measure can be used to compare the low- and high-dimensional representations and provide a quality measure for projections of categorical data. In this paper, we made qualitative statements about the effectiveness of dimensionality reduction methods for projecting categorical data. The next logical step is a quantitative evaluation of dimensionality methods, also considering the different types of representations and distance measures.
VII Conclusion

### v4

@@ 3 Interpreting the Categorical Data Map
In the following, we perform a case study on cluster and attribute analysis, using the Property Sales dataset [48] (see The Categorical Data Map: A Multidimensional Scaling-Based Approach) to show how to interpret emerging patterns for cluster, outlier, and similarity analysis. We chose this dataset because of its relative simplicity. However, it lacks the complexity of large categorical datasets, which we will address in an expert study (see Sec. 4).
@@
Cluster Analysis: There exist a total of $\Pi_{n\in\{1,\dots,|\mathcal{A}|\}}\,\left|\mathcal{C}_{n}\right|$ possible data items, given that all combinations of attributes are allowed, resulting in an exponential growth in the number of possible and unique data items. Hence, we can assume that there are dependencies and relationships among the categories contained in a dataset impacting their distribution. This means that groups of subsets that share a set of attributes should form perceivable structures (i.e., clusters) when projected using DR methods. Thus, our approach benefits from and leverages the sparsity of categorical data.
@@
For the Property Sales dataset, there are ten clusters (see The Categorical Data Map: A Multidimensional Scaling-Based Approach (1)). There is a symmetric split along the center of the projection. Given the size of this dataset, we can observe that the two attributes Purchaser Currently Living In and Location of Purchased Property dominate the appearance of the projection. The glyph sizes indicate that the categories Private Propriety and Central often occur together while {Private Propriety, Central, Condominium} is the largest unique subset. Thus, we can see that most private property is purchased in the central areas, and in this general group, the large majority are condominiums.
Dataset TW ($\uparrow$) CT ($\uparrow$) SC ($\uparrow$) NS ($\downarrow$) Avg. NH ($\uparrow$) Med. NH ($\uparrow$)
MDS+O
MDS+J
MCA
MDS+O
MDS+J
MCA
MDS+O
MDS+J
MCA
MDS+O
MDS+J
MCA
MDS+O
MDS+J
MCA
MDS+O
MDS+J
MCA
Audiology [11] .58 .89 .83 .64 .90 .89 .29 .77 .69 .17 .09 .81 .89 .92 .92 .98 .98 .98
Mushroom [55] .96 .97 .91 .92 .93 .97 .78 .77 .79 .08 .09 .64 .89 .90 .84 .90 .92 .87
Titanic [26] .86 .86 .76 .84 .84 .81 .76 .75 .59 .07 .07 .28 .68 .68 .63 .74 .75 .60
Cyber-Security [86] .84 .87 .79 .83 .86 .81 .87 .82 .68 .04 .06 .30 .56 .55 .54 .66 .62 .58
Property Sales [48] .91 .89 .73 .86 .85 .81 .70 .66 .46 .09 .10 .24 .65 .65 .51 .76 .76 .39
HCI Study (1) [67] .84 .77 .71 .81 .79 .73 .72 .69 .61 .07 .07 .18 .59 .58 .58 .53 .53 .52
HCI Study (2) [67] 1.0 1.0 1.0 1.0 1.0 1.0 .86 .85 .89 .03 .03 .11 .56 .56 .56 .57 .57 .57
Table 1: We compare projections of MDS using the Overlap coefficient (MDS+O) and the Jaccard distance (MDS+J) to MCA by applying them to seven real-world datasets. The MDS outperforms MCA for most datasets and quality metrics. In the case of the Audiology dataset with high category overlap, usually present in datasets with many attributes, we found that MDS combined with the Jaccard distance outperforms both alternatives.
@@
Attribute Analysis: By encoding the attribute values in the background, we enable users to analyze the distribution of subsets in the projection with respect to one or two attributes. For the Property Sales dataset, we found that the attribute Purchaser Currently Living In creates a clear and straight division between subsets (see The Categorical Data Map: A Multidimensional Scaling-Based Approach (2)). We can also see a second level of grouping by the Location of Purchased Property attribute forming a close to orthogonal split in the projection, which can be spotted with our visualizations (see The Categorical Data Map: A Multidimensional Scaling-Based Approach (3) and Fig. 5). Thus, Purchaser Currently Living In and Location of Purchased Property are the primary attributes. This finding is substantiated when checking the side panel entry of the attribute Property Type Purchased, which has three categories with low frequency. The appearance of the partitioning depends a lot on the selected attributes. When observing the layout enrichment, attributes present themselves on a spectrum from a few clearly separated groups to intermingled and highly fractured appearances. Property Type Purchased does not contribute to elements’ clustering (or cluster cohesion) since most groups contain subsets of the majority of its categories. Thus, the areas of the categories are disjointed, which reflects the fact that most property types are sold as both private and public property, as well as most of the geographic locations.
@@
Figure 6: Two visualizations of the Titanic dataset [26]. A split Euler diagram without the Age attribute (left) and an overlap reduced Parallel Sets visualization (right) with very thin ribbons. Both have drawbacks with a small dataset and do not scale with an increasing number of attributes.
Figure 7: Categorical Data Map visualizations of the Titanic dataset [26] using MDS and Overlap coefficient [77]. (1) The visualization shows six clusters and two outliers. The largest cluster is the subset of Adult, Male, Perished (at the bottom). The background encoding shows that the Survived and Sex attributes are relevant for this dataset, clearly separating the data items. For Sex, the separation is left and right. (2) For Survived, the separation is bottom-right/top-left. (3) The Age dimension also yields a separation, while (4) Class shows no clear structure.
4 Evaluation
@@
We qualitatively compare our Categorical Data Map to existing visualizations for categorical data and quantitatively compare our approach to MCA used by Broeksema et al. [19]. Additionally, we performed an expert study on two representative datasets with five data scientists.
4.1 Comparison to Euler Diagrams and Parallel Sets
@@
For categorical data, each data point has exactly one category for each attribute, while in Euler diagrams, the number of sets an element is included in is not restricted, i.e., it could be in less. Thus, to truthfully represent categorical data in Euler diagrams, there need to be $\Sigma_{a_{i}\in\mathcal{A}}\,|\mathcal{C}_{i}|$ sets, i.e., one set for each category of all attributes. Euler diagrams may require the selection of specific subsets of attributes and, therefore, are less suitable for exploratory data analysis. For highly intersecting sets, automatic layout methods might not create a single diagram [62]. We show an example of an automatically generated split Euler diagram for the Titanic dataset in Fig. 6 (left). The attribute Age was removed to reduce the diagram’s complexity. The Titanic dataset requires ten sets. However, even with eight sets, the visualization is disjointed. Parallel Sets are alternative categorical sets visualization, combining principles from stacked bars and parallel coordinate plots [15, 49]. Fig. 6 (right) shows the Titanic dataset in a Parallel Sets visualization, where the readability is improved through overlap reduction. Small subsets are represented as very thin ribbons on the lowest level, which can be hard to perceive. Visualizing the Mushroom dataset with classical Parallel Sets is not visually feasible since it will have 22 ribbon layers and 8123 subsets on the lowest level (see supplementary material). Alsakran et al.  [3] addressed this issue by only visualizing 2-dimensional subsets in a modified Parallel Sets visualization. However, the relation between 2-dimensional subsets is lost. Thus, we argue that Euler diagrams and Parallel Sets, as examples of established visualizations for categorical data, do not scale with an increasing number of attributes.
4.2 Quantitative Evaluation of Projection Quality
We use five quality metrics commonly used in related work for DR to evaluate and compare the quality of our categorical data projections [32]. The result of comparing MDS with Overlap coefficient (MDS+O) and Jaccard distance (MDS+J) to MCA are shown in Tab. 1. We briefly describe each metric below and use them to compare our MDS-based method to MCA using seven real-world categorical datasets.
Trustworthiness (TW) [82] quantifies the proportion of points that remain close in the lower-dimensional representation to assess how accurately local patterns in the projection represent the data patterns. This is linked to the occurrence of "false neighbors" in the protection.
Continuity (CT) [82] measures the ratio of points in the projection that remain close in the original space. This is related to the "missing neighbors" of a projected point.
Normalized Stress (NS) [41] quantifies how well the distances between pairs of points are preserved when mapping from the original space to the projected space. This measure should be as low as possible.
Shepard Diagram Correlation (SC) [31] measures the rank correlation of all distances of the original and the projected space, assessing the quality of distance preservation globally using Spearman’s $\rho$ [74].
Neighborhood Hit (NH) [64] measures the proportion of a point’s neighbors in the projection space that share the same label as the point itself, averaged across all points in its neighborhood. This metric is related to the separation of labeled data in the projection. In our case, we evaluate every attribute as a set of labels. Thus, we calculate the mean and median values of NH across all attributes of a dataset.
TW, CT, NH require a parameter $k$ defining a neighborhood size. We set $k=7$, a commonly used value [31]. We found that our approach generally outperforms MCA quantitatively. Additional measurements of MDS with other distance measures, detailed descriptions of the quality metrics, and datasets are provided in the supplementary material.
4.3 Qualitative Expert User Study
@@
To evaluate the Categorical Data Map we performed a paired analytics study [43]. We conducted an expert study with five data scientists, E1–E5, with varying backgrounds. All participants were Ph.D. candidates and students. All were male, and the age range was 25 to 30 years. All experts had experience in the area of information visualization and visual analytics. During the study, we asked the experts to verbalize their thought process to capture it. The following studies are set up using MDS projections of the Mushroom and Titanic dataset using the Overlap coefficient (MDS+0). Tab. 1 shows that these projections are higher quality than MCA-based ones regarding most quality metrics.
@@
All trials followed a predefined structure and took between 43 and 57 minutes. The study was conducted in German. The study started with an introduction to the Categorical Data Map using the Property Sales dataset by Hassan et al. [36] shown in The Categorical Data Map: A Multidimensional Scaling-Based Approach and included a description of the square area glyph, layout enrichment, and interactions to introduce the expert to the prototype. After the introduction, the experts had the opportunity to ask questions regarding our approach.
@@
Titanic Dataset: The experts had to analyze the Titanic dataset [26] using the Categorical Data Map shown in Fig. 7. E1–E5 were able to locate the largest subset {Male, Perished, Adult, Crew} by looking at the visualization without any additional interaction (Fig. 7 (1)). E1–E5 used Lasso selection to find and validate that the largest subset regarding three attributes is {Male, Perished, Adult} (Fig. 7 (2)). Additionally, E1–E5 were able to find six clusters and two outliers. E1, E3, and E4 found that the outliers represent the subsets defined by the categories {Perished, Child, Crew} (Fig. 7 (3)). E1, E3, and E5 commented on the high number of perished males and the large number of casualties among the {Male, Crew}. E1–E5 used the layout enrichment to navigate and reason about the location of subsets, including the Class attribute (Fig. 7 (4)). E2 commented on the close to orthogonal split in the projection between Sex and Survived shown in Fig. 7 (1) and (2).
@@
Figure 8: Categorical Data Map visualizations of the Mushroom dataset [55] using the MDS and Overlap coefficient. (1) Two poisonous mushrooms very similar to edible mushrooms. (2) Comparing class and odor reveals that the poisonous outlier has a pungent odor. Continued analysis reveals that mushrooms with an unpleasant smell are poisonous. (3) After the selection of a cluster, the ring-type is identified as a defining characteristic for the cohesion of visible clusters and is used as a property for the classification of mushrooms. (4) Selecting two poisonous clusters, reveals that the vast amount of poisonous mushrooms are silky at the stalk-surface-below-ring, while there exist very few silky mushrooms that are edible.
@@
Mushroom Dataset: E1–E5 had the opportunity to perform an open exploration task and were only given the information that the dataset is about mushrooms and that the class attribute indicates their poisonousness. The glyph was replaced with a simple black dot to reduce the visual complexity. E1–E5 perceived five clusters right at the outset. E2 and E3 used the Lasso selection together with layout enrichment to determine differentiating categories for cluster separation, e.g., evanescent, large, and pendent for the ring-type attribute (Fig. 8 (3)). E1–E5 found the poisonous outliers nested in the group representing edible mushrooms (Fig. 8 (1)) being poisonous mushrooms very similar to edible ones. E1–E5 found the general rule that mushrooms with an fishy foul, musty, spicy or other unpleasant smells indicate a poisonous mushroom (Fig. 8 (2)). During the open exploration task, E1, E2, and E3 found the rule without additional information. E4 and E5 needed help to find the class and odor combination. However, E4 and E5 could deduce the rule by only interpreting the plot. E3, quickest in exploring the dataset, found that stalk-surface-below-ring is silky for the majority of poisonous mushrooms and the stalk-surface-below-ring is mostly smooth for edible ones (Fig. 8 (4)).
@@
General Comments: Before concluding the study, the participants were asked to comment on their preferences for the available glyph designs. E1, E2, and E4 preferred a circular glyph design (Fig. 1 (c) and (d)) over a square design. E3 and E5 preferred square glyph designs (Fig. 1 (a) and (b)). E1 and E3 found that the area-based glyphs are inferior to the alternative designs for reading off precise subset sizes. E1 mentioned as a drawback that the glyphs are not rotation invariant. E1 commented that the layout enrichment is very useful for navigation and orientation and helps to perceive the impact on category groups. However, E1 also noted that the layout enrichment does not reflect the ratio of data items with a given category. E3 mentioned a general preference for the map metaphor by being helpful for orientation among different subsets. E2 mentioned potential scalability issues with the glyph for large datasets, e.g., for a high number of attributes, and proposed semantic zoom as a potential option. E1–E5 commented that ordering attributes according to their fracturedness was understandable and useful. During the general questions at the end, E2–E4 freely explored plots created with other distance measures and DR methods. E3 commented that the result of MCA-based plots was hard to interpret, noticing the disjointed layout enrichment and thus having larger fracturedness. E1 mentioned issues with the encoding of categories, such as the category North not being located north of the plot or the category brown not having the color brown, and suggested being able to select the color of a category manually.
5 Discussion and Future Work
@@
In this section, we discuss the lessons-learned, reflect on the design decisions, and discuss computational complexity and future work.
@@
Visualizing Attributes and Categories: We initially used circular glyphs as shown in Fig. 1 (c) and (d), which had the benefit of using the available space effectively since overlap minimization relative to the radius is straightforward to implement. The subset size encoding by the arc around the circle enables finer-grained distinction of sizes since it offers more space. However, during the design phase, users misinterpreted the circle segments as pie charts, a common method for displaying categorical data. Thus, we decided to circumvent this common misconception by using square-based representation for the categorical subsets. However, three out of five experts preferred a circular glyph design.
@@
Encoding of Subset Sizes: We evaluated four different visual encodings for the size of a categorical data subset (see Fig. 1). The area-based glyph makes it easier to perceive subset sizes at a glance, and thus, a user can spot the distribution of the dataset directly. Still, it suffers from overlap, especially for tight clusters. Thus, there is a benefit to applying methods to reduce overlap. We are able to mitigate the overlap problem with the force-directed overlap reduction largely. Simplifying the representation of a dot requires less space, but the assessment of subset sizes requires interaction. It is possible to remove the subset size information altogether. However, this may limit analysis tasks where the subset sizes it not important, e.g., the Mushroom dataset. All glyph designs benefit from a mouse-over mechanism that moves the currently selected glyph to the top so that all attributes can be observed.
@@
Encoding of Attributes Into the Background: Fig. 7 shows that encoding an attribute into the visualization gives insight into the topology of the projection. We could also show the benefit of encoding multiple attributes into the background to allow for a more complex representation of the topology. We found that the number of categories of an attribute weakly influences the fracturedness of an attribute. However, the main factor is the number of subsets containing the attribute, i.e., an attribute with two categories and an occurrence roughly equal among all subsets will yield a low fracturedness for that attribute. With increased imbalance between the categories, the fracturedness may increase if other more balanced attributes are present.
@@
We discussed the use of weighted Voronoi diagrams [5] to better reflect the subset size in the background encoding. The use of a weighted Voronoi diagram will conflict with local cluster patterns; more specifically, for imbalanced datasets, the area of one Voronoi cell extends below the point of its neighbors, requiring restrictions on the range weights. This behavior makes the layout enrichment hard to interpret since points are placed inside or close to an area representing a category they do not belong to. For datasets with only unique entries, the weight Voronoi diagram will be identical to the regular Voronoi diagram. To organize subsets, we also considered Voronoi Treemaps [10]. However, Voronoi Treemaps require a hierarchical structure, just like regular Tree Maps [71] and, thus, cannot be applied to categorical data without additional information to derive a hierarchy of attributes.
@@
Computational Complexity: The number of data records $n$ poses potential limitations. The time complexity of projecting data is determined by the DR methods. However, since categorical data sets are sparse, as discussed in Sec. 3, the number of projected subsets is significantly lower than that of data records. The Voronoi diagram calculation and the corresponding Delaunay triangulation are both in $O(n\leavevmode\nobreak\ log(n))$ [9]. The time complexity of calculating the fracturedness measures depends on the number of vertices and edges of the Delaunay triangulation, which will have $n$ vertices and $3n-3-h$ edges, where $h$ is the number of vertices on the convex hull. The time complexity of calculating edge-based fracturedness is based on enumerating all edges of the Delaunay triangulation and has a time complexity of $O(|E|)$. The time complexity of calculating component-based fracturedness is dependent on the algorithm for determining the number of components. We use a depth-first search-based approach with time complexity of $O(|V|+|E|)$. Thus, the dimensionality reduction method employed poses the highest contribution to the time complexity, $O(n^{3})$ for MDS.
Future Work: We found that Voronoi cells can overrepresent the amount of data associated with a specific category. Thus, there is a need for a new layout enrichment method following these constraints: (1) The global area associated with one category should be relative to the occurrence in the dataset (data-ink ratio), (2) the extent of individual category areas should remain close to their projected data point positions, (3) where meaningful (e.g., among clusters), the layout enrichment should visually enclose the data points with the same category if the data-ink ratio allows. The expert study showed that the color assignment for foreground and background colors could be improved. We suggest assigning attributes to a few sets of colors based on an exploration phase. Later in the analysis, we require one color set for the attribute used in the background, one for the foreground using a distinctive palette, and one for the attribute under focus by the user. All the other attributes would be assigned a neutral color (e.g., grey). In this paper, we studied the use of MDS for categorical data analysis. However, following the approach of encoding categorical data into distances, other DR methods could be used (e.g., t-SNE [80] or UMAP [56]). These can be evaluated and compared quantitatively, following the evaluation presented in this paper. We think that the concept of fracturedness can be transferred to high-dimensional space when analyzing categorical data. Such a measure can be used to compare the low- and high-dimensional representations and provide a quality measure for projections of categorical data.
6 Conclusion